# Optimizing Distributed Reinforcement Learning with Reactor Model and Lingua Franca

### 07 Dec 2023 v1

@@ 5.1 Number of Actors
Fig. 4 illustrates the mean overhead of broadcasting and gathering a 10MB object across different numbers of actors in a distributed RL setup. Two frameworks are compared: Ray and Lingua Franca (LF). The x-axis represents the number of actors, which are 2, 4, 8, and 16, while the y-axis shows the overhead in milliseconds. Both frameworks exhibit an increase in overhead as the number of actors grows, but Ray consistently has a higher overhead than LF. For instance, with 16 actors, Ray’s overhead is close to 20 milliseconds, whereas LF’s is just above 5 milliseconds. The error bars suggest variability in the measurements, with Ray showing larger variance as the number of actors increases.
5.2 Object Size
Fig. 4 presents a comparison of the mean overhead for broadcasting and gathering operations on objects of varying sizes using 16 actors, between Ray and Lingua Franca (LF) frameworks, including a 99% confidence interval (CI). The x-axis displays the object size in megabytes (MB), ranging from 0 to 500 MB, while the y-axis indicates the overhead in milliseconds. From the graph, we can observe that as the object size increases, the overhead for both Ray and LF also increases. However, Ray’s overhead grows at a higher rate than LF’s. For instance, with the largest object size of 500 MB, Ray’s overhead approaches 800 milliseconds, whereas LF’s overhead is about half of that, around 400 milliseconds.
Both figures 4 and 4 support the hypothesis that the reactor model employed by LF can outperform traditional actor models like Ray, particularly in scenarios with a greater number of actors and larger object size, due to optimizations such as reduced synchronization work and efficient multithreading with No GIL Python.
Figure 5: Simulation Throughput of Ray and LF in Various Gym Environments.
Figure 5: Simulation Throughput of Ray and LF in Various Gym Environments.
@@ 5.2 Object Size
Figure 6: Simulation Throughput of Ray and LF in Various Atari Environments.
5.3 Open AI Gym Environments
In Fig. 6, we see that LF outperforms Ray by 1.21x on average, with a particularly significant lead in the Blackjack environment. This suggests that LF is more efficient, especially in situations where there is a lower CPU demand for action updates within the environment. Considering the CPU architecture in m5x8.xlarge, which has a mix of 8 performance cores and 16 efficiency cores, the observed performance is not ideal due to the differing speeds of the cores and potential thermal throttling. The efficiency cores are slower and can bottleneck the system as the benchmark is limited by the slowest thread. Moreover, when multiple cores are active, the CPU may not run at its highest frequency due to thermal constraints. To optimize performance and achieve linear scaling, it is recommended to utilize only the 8 performance cores and disabled. This can be done by setting CPU affinity for performance cores, ensuring that CPU-bound tasks, which require synchronization, do not get restricted by the slower efficiency cores. By running tasks only on the performance cores, one can avoid the inconsistencies that come with hybrid architectures and get a more linear scaling with the number of actors or threads in use. It’s important to note that vectorized environments are asynchronous and do not parallelize inference of policy. Therefore, it is not included in our benchmarks.
5.4 Atari Environments
@@
Figure 7: Dataflow Graph of Parallel Q-learning
Figure 8: Synchronized Parallel Q-learning with Different Batch Sizes from a Replay Buffer
5.5 Synchronized Parallel Q-learning
@@
Deep Q-Networks (DQNs)  mnih2013playing are an advancement in reinforcement learning that utilize deep neural networks to estimate Q-values. The Q-values are predictions of the expected rewards for taking certain actions given particular states in an environment. DQNs extend the capabilities of traditional Q-learning by handling larger state and action spaces, which are common in complex problems. These networks can scale up effectively with more data or increased model complexity; therefore, DQN usually utilizes GPU for gradient updates. However, since AWS EC2 m5.8xlarge doesn’t include a GPU, benchmarking was conducted on a workstation with an Intel i9-13950HX CPU @ 2.20GHz featuring 32 vCPUs, NVIDIA RTX4090, and 32 GiB of RAM. The system also runs on Ubuntu 20.04 and uses Python 3.9.10, with numpy 1.22.3, torch 1.9.0, and gym 0.20.0.
@@
DQN has been implemented in a synchronized parallel Q-learning setup within a BlackJack environment silver2013playing . The dataflow graph is shown in fig. 7. The network is trained to take a blackjack hand as input and output scores for each of the possible actions in the game, which represent the expected rewards of taking those actions. The computational tasks are distributed with the rollout and replay buffer being executed on a CPU, while the DQN Reactor uses a GPU. This setup takes advantage of GPUs for complex matrix operations and multicore CPUs for sequential decision-making simulations in Open AI Gym environments.
@@
The benchmark results, as seen in fig. 8, demonstrate performance improvements using LF over Ray. In tests with a 500 sample batch size from the replay buffer, the average training time decreased by 31.2%. The data also shows that while Ray’s training time increases with larger mini-batch sizes, LF’s performance remains stable, indicating LF’s ability to handle larger batches efficiently without significantly increasing resource utilization.
Figure 9: TrafficJunction4-v0 environment. The goal is to pass a four-way junction without any collisions.
Figure 9: TrafficJunction4-v0 environment. The goal is to pass a four-way junction without any collisions.
Figure 10: Multi-Agent Reinforcement Learning
5.6 Multi-Agent RL Inference Comparison
@@
In Multi-Agent RL (MARL), multiple agents operate in a common environment, where each of them tries to optimize its own return by interacting with the environment and other agents zhang2021multi (See fig. 10). In centralized MARL, a central controller aggregates information across the agents, including joint actions, rewards, and observations, and policies across different agents. In decentralized MARL, which is more common in cooperative situations, each agent makes decisions based on its local observations oroojlooy2023review .
@@
We validate LF in a MARL setting with ma-gym magym , a MARL library based on OpenAI Gym brockman2016openai . We use the TrafficJunction4-v0 environment, in which four agents are trying to pass a crossroad without crashing (see Fig. 10). This is a decentralized setting where each agent gets it’s own local observation. See in fig. 12, LF requires significantly less inference time than Ray (less than 1/2 the fitted line’s slope). It is also noteworthy that as the number of agents increases as shown in fig. 12, LF’s inference time scales better (again, less than 1/2 the rate of increase than that of Ray). In TrafficJunction environments with 10 Agents, LF achieves a 5.12x speed up compared to Ray.
Figure 11: Inference Time Between Ray and LF over Episode Counts
Figure 11: Inference Time Between Ray and LF over Episode Counts
@@
Figure 12: Inference Time Comparison between Ray and LF across Various Numbers of Agents
6 Discussion and Conclusion
@@
In this study, we introduced LF, a novel distributed RL framework based on the Reactor model, which outperforms widely used frameworks like Ray in handling training, serving, and simulation tasks in RL. It achieves this through reducing work needed for synchronization using Reactor model and lock-free data structure. Our empirical evaluations demonstrate LF’s superior performance: a 1.21x and 11.62x faster sample generation in OpenAI Gym and Atari environments, a 31.2% reduction in average training time for synchronized parallel Q-learning, and a 5.12x acceleration in multi-agent RL inference. We aim to incorporate the optimizations into LF’s federated execution, enabling efficient distributed training and serving across clusters. With this work, we make important contributions to the field of distributed RL.

### 02 Feb 2024 v2

@@ 5.2. Number of Actors
Fig. 4 illustrates the mean overhead of broadcasting and gathering a 10MB object across different numbers of actors in a parameter server setup. Two frameworks are compared: Ray and LF. The x-axis represents the number of actors, which are 2, 4, 8, and 16, while the y-axis shows the overhead in milliseconds. Both frameworks exhibit an increase in overhead as the number of actors grows, but Ray consistently has a higher overhead than LF. For instance, with 16 actors, Ray’s overhead is close to 20 milliseconds, whereas LF’s is just above 5 milliseconds.
Figure 6. Simulation Throughput of Ray and LF with 16 actors in Various Gym Environments.
@@
Figure 7. Simulation Throughput of Ray and LF with 16 actors in Various Atari Environments.
5.3. Object Size
Fig. 5 presents a comparison of the mean overhead for broadcasting and gathering operations on objects of varying sizes using 16 actors, between Ray and LF frameworks, including a 99% confidence interval (CI). The x-axis displays the object size in megabytes (MB), ranging from 0 to 500 MB, while the y-axis indicates the overhead in milliseconds. From the graph, we can observe that as the object size increases, the overhead for both Ray and LF also increases. However, Ray’s overhead grows at a higher rate than LF’s. For instance, with the largest object size of 500 MB, Ray’s overhead approaches 800 milliseconds, whereas LF’s overhead is about half of that, around 400 milliseconds.
Both Fig. 4 and Fig.5 support the hypothesis that the reactor model employed by LF can outperform traditional actor models like Ray, particularly in scenarios with a greater number of actors and larger object size, due to optimizations such as reduced synchronization work and efficient multithreading with No GIL Python.
5.4. Open AI Gym Environments
In Fig. 6, we see that LF outperforms Ray in terms of simulation throughput for Open AI Gym Environments by 1.21x on average, with a particularly significant lead in the Blackjack environment. This suggests that LF is more efficient, especially in situations where there is a lower CPU demand for action updates within the environment. It’s important to note that vectorized environments are asynchronous and do not parallelize inference of policy. Therefore, it is not included in our benchmarks.
5.5. Atari Environments
@@
Figure 8. Dataflow Graph of Parallel Q-learning
5.6. Synchronized Parallel Q-learning
@@
Deep Q-Networks (DQNs)  (Mnih et al., 2013) are an advancement in reinforcement learning that utilize deep neural networks to estimate Q-values. The Q-values are predictions of the expected discounted returns after taking certain actions given particular states in an environment. DQNs extend the capabilities of traditional Q-learning by handling larger state and action spaces, which are common in complex problems. These networks can scale up effectively with more data or increased model complexity; therefore, DQN usually utilizes GPUs for gradient updates. However, since AWS EC2 m5.8xlarge does not include a GPU, benchmarking was conducted on a workstation with an Intel i9-13950HX CPU @ 2.20GHz featuring 32 vCPUs, NVIDIA RTX4090, and 32 GiB of RAM. The system also runs on Ubuntu 20.04 and uses Python 3.9.10, with numpy 1.22.3, torch 1.9.0, and gym 0.20.0.
@@
DQN has been implemented in a synchronized parallel Q-learning setup within a BlackJack environment (Silver et al., 2013). The dataflow graph is shown in Fig. 8. The network is trained to take a blackjack hand as input and output scores for each of the possible actions in the game, which represent the expected rewards of taking those actions. The computational tasks are distributed with the rollout and replay buffer being executed on a CPU, while the DQN Reactor uses a GPU. This setup takes advantage of GPUs for complex matrix operations and multicore CPUs for sequential decision-making simulations in Open AI Gym environments.
Figure 9. Synchronized Parallel Q-learning with Different Batch Sizes from a Replay Buffer
@@
The benchmark results, as seen in Fig. 9, demonstrate performance improvements using LF over Ray. In tests with a 500 sample batch size from the replay buffer, the average training time decreased by 31.2%. The data also shows that while Ray’s training time increases with larger mini-batch sizes, LF’s performance remains stable, indicating its ability to handle larger batches efficiently.
5.7. Multi-Agent RL Inference Comparison
@@
In Multi-Agent RL (MARL), multiple agents operate in a common environment, where each of them tries to optimize its own return by interacting with the environment and other agents (Zhang et al., 2021). In centralized MARL, a central controller aggregates information across the agents, including joint actions, rewards, and observations, and policies across different agents. In decentralized MARL, which is more common in cooperative situations, each agent makes decisions based on its local observations (Oroojlooy and Hajinezhad, 2023).
Figure 10. Inference Time Between Ray and LF over Episode Counts
@@
We validate LF in a MARL setting with ma-gym (Koul, 2019), a MARL library based on OpenAI Gym (Brockman et al., 2016). We use the TrafficJunction4-v0 environment, in which four agents are trying to pass a crossroad without crashing. This is a decentralized setting where each agent gets its own local observation. As shown in Fig. 10, LF requires significantly less inference time than Ray (less than 1/2 the fitted line’s slope). It is also noteworthy that as the number of agents increases as shown in Fig. 11, LF’s inference time scales better (again, less than 1/2 the rate of increase than that of Ray). In TrafficJunction environments with 10 agents, LF achieves a 5.12x speed-up compared to Ray.
@@
Figure 11. Inference Time Comparison between Ray and LF across Various Numbers of Agents
6. Discussion
@@
We demonstrate that LF outperforms widely used frameworks like Ray in handling training, serving, and simulation tasks in RL. We achieve this by reducing the work needed for synchronization using the reactor model and decreasing the I/O overhead through optimizing the coordination of Python worker threads. Our empirical evaluations demonstrate LF’s superior performance: a 1.21x and 11.62x higher simulation throughput in OpenAI Gym and Atari environments, a 31.2% reduction in average training time for synchronized parallel Q-learning, and a 5.12x acceleration in multi-agent RL inference. We aim to incorporate the optimizations for single-node described in this work into LF’s federated execution, enabling efficient distributed training and serving across nodes. We also plan to delve deeper into the potential applications of our optimizations in deploying deep RL on embedded systems, and will compare it with frameworks such as the Robot Operations System, a middleware commonly used in robotics.